Steam flows through a pipe at the rate of 5kg/s. The pressure nd temperature are 12 bar nd 300 degree celsius respectively . If 2000J of heat is lost to the surrounding at constant pressure . Find the final condition of the steam.
1
Expert's answer
2019-04-05T10:22:01-0400

If 2000J is the heat lost per second, then Q = νcp(T1 – T2), respectively T2 = T1 – Qμ/4Rm = 573 – 36/(4*8.31*5) = 572.8 (K).